# Chinese SimpleQA: A Chinese Factuality Evaluation for Large Language Models

### v1

@@ 1 Introduction
However, the SimpleQA benchmark primarily targets the English language, resulting in a limited understanding of LLMs’ capabilities in other languages. Moreover, inspired by several recent Chinese LLM benchmarks (e.g., C-Eval (Huang et al., 2023), CMMLU (Li et al., 2023b)), to evaluate the factuality abilities of LLMs in Chinese, we present the Chinese SimpleQA benchmark [^1] https://openstellarteam.github.io/ChineseSimpleQA/, which consists of 3000 high-quality questions spanning 6 major topics, ranging from humanities to science and engineering, as shown in Figure 1. Specifically, the distinct main features of Chinese SimpleQA are as follows:
@@ 3.1 Baseline Models
o1-preview 63.8 8.1 24.0 72.7 67.9 45.7 69.8 72.4 65.0 73.5 72.3
Doubao-pro-32k 61.9 5.8 27.8 69.1 65.3 61.8 69.3 69.0 56.1 64.2 70.4
GLM-4-Plus 58.7 2.8 33.9 63.4 60.9 56.5 64.1 64.9 50.7 66.6 62.8
GPT-4o 59.3 1.3 39.3 60.1 59.7 39.4 64.0 65.1 53.3 68.6 62.0
Qwen-Max 54.1 5.8 34.6 61.0 57.4 47.8 59.9 63.5 49.9 61.2 59.3
Gemini-1.5-pro 54.4 2.9 37.6 59.1 56.7 41.4 59.1 60.8 52.2 56.3 64.3
DeepSeek-V2.5 54.1 1.8 40.0 57.5 55.7 50.4 57.6 58.8 50.1 59.4 56.9
Claude-3.5-Sonnet 46.2 23.3 26.4 63.6 53.5 28.7 61.3 60.4 42.2 59.8 57.7
Yi-Large 47.3 12.2 36.3 56.6 51.5 41.1 56.5 55.1 41.7 57.6 53.8
moonshot-v1-8k 48.7 1.0 45.9 51.5 50.1 49.8 54.1 56.8 41.4 53.0 46.6
GPT-4-turbo 45.6 14.3 40.2 53.1 49.1 24.2 55.2 58.9 43.9 52.5 50.8
GPT-4 45.4 3.7 46.2 49.6 47.4 25.2 54.0 52.8 41.8 52.8 50.6
Baichuan3-turbo 45.2 4.8 45.8 49.6 47.3 32.3 52.5 54.0 35.4 54.6 50.9
o1-mini 39.5 20.5 39.9 49.7 44.1 21.3 49.2 55.9 33.8 48.8 46.8
Doubao-lite-4k 36.7 26.7 32.1 53.3 43.4 40.2 44.8 51.0 31.1 41.4 50.4
@@
Qwen2.5-72B 48.4 1.8 44.5 52.1 50.2 36.3 56.1 57.9 37.1 53.3 56.4
Qwen2.5-32B 38.8 5.6 50.1 43.6 41.1 33.7 45.8 48.7 27.3 44.7 44.9
Qwen2.5-14B 35.4 4.2 55.0 39.2 37.2 30.2 41.8 46.1 24.1 38.8 41.0
Qwen2.5-7B 26.6 4.6 63.5 29.5 27.9 20.1 32.7 33.8 18.0 28.6 32.0
Qwen2.5-3B 16.2 7.6 71.0 18.6 17.3 13.4 17.9 26.1 9.3 15.6 20.8
Qwen2.5-1.5B 11.1 9.3 74.3 13.1 12.0 11.0 11.3 18.7 6.7 12.2 12.9
GLM4-9B 25.9 8.3 61.6 29.6 27.6 28.8 32.1 32.0 17.6 28.9 27.8
ChatGLM3-6B 11.2 9.4 75.2 12.9 12.0 12.1 13.8 12.4 8.8 13.4 11.8
@@
DeepSeek-V2-Lite-Chat 33.7 8.7 53.5 38.6 36.0 35.3 38.5 41.7 32.2 37.5 31.2
@@ 3.2 Main Results
  • Small LLMs usually lead to higher scores on “not attempted (NA)”. The NA scores for o1-mini, InternLM2.5-1.8B are 20.5 and 9.3, respectively, which are larger than the scores of corresponding larger LLMs a lot (o1-preview with 8.1, Qwen2.5-72B with 1.8).
@@ 5 Conclusion
In this paper, to evaluate the factuality abilities of existing LLMs, we propose the first Chinese short-form factuality benchmark (i.e., Chinese SimpleQA), which includes 6 main topics and 99 subtopics. Besides, Chinese SimpleQA mainly has five important features (i.e., Chinese, diverse, high-quality, static, and easy-to-evaluate). Based on Chinese SimpleQA, we comprehensively evaluate the performance of existing 40+ LLMs on factuality and provide detailed analysis to demonstrate the advantage and necessity of our Chinese SimpleQA. In the future, we will investigate to improve the LLMs’ factuality and explore to extend Chinese SimpleQA to multilingual and multimodal settings.

### v2

@@ 1 Introduction
However, the SimpleQA benchmark primarily targets the English language, resulting in a limited understanding of LLMs’ capabilities in other languages. Moreover, inspired by several recent Chinese LLM benchmarks (e.g., C-Eval (Huang et al., 2023), CMMLU (Li et al., 2023b)), to evaluate the factuality abilities of LLMs in Chinese, we present the Chinese SimpleQA benchmark [^1] https://openstellarteam.github.io/ChineseSimpleQA/, which consists of 3000 high-quality questions spanning 6 major topics, ranging from humanities to science and engineering, as shown in Figure 1. Specifically, the distinct main features of our proposed Chinese SimpleQA dataset are as follows:
@@ 3.1 Baseline Models
o1-preview 63.8 12.2 24.0 72.7 67.9 45.7 69.8 72.4 65.0 73.5 72.3
Doubao-pro-32k 61.9 10.3 27.8 69.1 65.3 61.8 69.3 69.0 56.1 64.2 70.4
GLM-4-Plus 58.7 7.4 33.9 63.4 60.9 56.5 64.1 64.9 50.7 66.6 62.8
GPT-4o 59.3 1.4 39.3 60.1 59.7 39.4 64.0 65.1 53.3 68.6 62.0
Qwen-Max 54.1 11.3 34.6 61.0 57.4 47.8 59.9 63.5 49.9 61.2 59.3
Gemini-1.5-pro 54.4 8.0 37.6 59.1 56.7 41.4 59.1 60.8 52.2 56.3 64.3
DeepSeek-V2.5 54.1 5.9 40.0 57.5 55.7 50.4 57.6 58.8 50.1 59.4 56.9
Claude-3.5-Sonnet 46.2 27.4 26.4 63.6 53.5 28.7 61.3 60.4 42.2 59.8 57.7
Yi-Large 47.3 16.4 36.3 56.6 51.5 41.1 56.5 55.1 41.7 57.6 53.8
moonshot-v1-8k 48.7 5.4 45.9 51.5 50.1 49.8 54.1 56.8 41.4 53.0 46.6
GPT-4-turbo 45.6 14.2 40.2 53.1 49.1 24.2 55.2 58.9 43.9 52.5 50.8
GPT-4 45.4 8.4 46.2 49.6 47.4 25.2 54.0 52.8 41.8 52.8 50.6
Baichuan3-turbo 45.2 9.0 45.8 49.6 47.3 32.3 52.5 54.0 35.4 54.6 50.9
o1-mini 39.5 20.6 39.9 49.7 44.1 21.3 49.2 55.9 33.8 48.8 46.8
Doubao-lite-4k 36.7 31.2 32.1 53.3 43.4 40.2 44.8 51.0 31.1 41.4 50.4
@@
Qwen2.5-72B 48.4 7.1 44.5 52.1 50.2 36.3 56.1 57.9 37.1 53.3 56.4
Qwen2.5-32B 38.8 11.1 50.1 43.6 41.1 33.7 45.8 48.7 27.3 44.7 44.9
Qwen2.5-14B 35.4 9.6 55.0 39.2 37.2 30.2 41.8 46.1 24.1 38.8 41.0
Qwen2.5-7B 26.6 9.9 63.5 29.5 27.9 20.1 32.7 33.8 18.0 28.6 32.0
Qwen2.5-3B 16.2 12.8 71.0 18.6 17.3 13.4 17.9 26.1 9.3 15.6 20.8
Qwen2.5-1.5B 11.1 14.6 74.3 13.1 12.0 11.0 11.3 18.7 6.7 12.2 12.9
GLM4-9B 25.9 12.5 61.6 29.6 27.6 28.8 32.1 32.0 17.6 28.9 27.8
ChatGLM3-6B 11.2 13.6 75.2 12.9 12.0 12.1 13.8 12.4 8.8 13.4 11.8
@@
DeepSeek-V2-Lite-Chat 33.7 12.8 53.5 38.6 36.0 35.3 38.5 41.7 32.2 37.5 31.2
@@ 3.2 Main Results
  • Small LLMs usually lead to higher scores on “not attempted (NA)”. The NA scores for o1-mini, InternLM2.5-1.8B are 20.5 and 31.2, respectively, which are larger than the scores of corresponding larger LLMs a lot (o1-preview with 12.2, InternLM2.5-20B with 7.7).
@@ 5 Conclusion
In this paper, to evaluate the factuality abilities of existing LLMs, we propose the first Chinese short-form factuality benchmark (i.e., Chinese SimpleQA), which mainly has five important features (i.e., Chinese, diverse, high-quality, static, and easy-to-evaluate). Based on Chinese SimpleQA, we comprehensively evaluate the performance of existing 40+ LLMs on factuality and provide detailed analysis to demonstrate the advantage and necessity of our Chinese SimpleQA. In the future, we will investigate how to improve the LLMs’ factuality and explore how to extend Chinese SimpleQA to multilingual, multimodal, and domain-specific (e.g., code, e-commerce) settings.